# When Does Multi-Task Learning Fail? Quantifying Data Imbalance and Task Independence in Metal Alloy Property Prediction

### v1

@@ Abstract
Multi-task learning (MTL) has emerged as a promising approach in materials informatics, with the assumption that related material properties share underlying physical principles that can be leveraged to improve prediction accuracy. In this study, we critically examine this assumption by applying MTL to simultaneously predict three properties of metal alloys: electrical resistivity, Vickers hardness, and amorphous-forming ability. Using a union dataset of 54,028 samples from the AI-Hub Korea platform, we compare three approaches: independent single-task models, standard MTL with hard parameter sharing, and structured MTL with learned task relation graphs. Contrary to expectations, our results reveal a striking dichotomy: MTL significantly degrades regression performance (resistivity $R^{2}$: 0.897 $\rightarrow$ 0.844, $p<0.01$; hardness $R^{2}$: 0.832 $\rightarrow$ 0.694, $p<0.01$) while significantly improving classification performance (amorphous F1: 0.703 $\rightarrow$ 0.744, $p<0.05$; recall improved by 17%). Analysis of the learned task relation graphs reveals near-zero inter-task weights ($\sim$0.006), indicating that these alloy properties are fundamentally independent rather than coupled. We attribute the regression degradation to negative transfer caused by severe data imbalance (52,388 resistivity samples vs. 800 hardness samples), where the dominant task overwhelms minority tasks during joint optimization. These findings provide empirical guidelines for practitioners: use independent models for precise property regression, but leverage MTL for classification tasks where improved recall is critical for candidate screening. Our work highlights the importance of validating MTL assumptions before deployment in materials discovery pipelines.
@@
[…]
[1]organization=Department of Electrical and Computer Engineering, Korea University, city=Seoul, postalcode=02841, country=Republic of Korea
@@ 2.2 Negative Transfer and Task Interference
Chen et al. [chen2018gradnorm] proposed GradNorm to dynamically balance gradient magnitudes across tasks, partially mitigating imbalance effects. Kendall et al. [kendall2018multi] introduced uncertainty-based task weighting. However, these methods address symptoms rather than the fundamental question of task compatibility.
@@ 2.4 MTL for Metal Alloys
Debnath et al. [debnath2025overcoming] used MTL to predict yield strength and elongation of HEAs, finding modest improvements. Pei et al. [pei2020machine] predicted hardness and phase stability using composition features. However, these studies typically use small, curated datasets where all samples have all properties labeled—a scenario quite different from the realistic setting where different properties are measured on different sample sets.
Our work addresses this realistic scenario using a union dataset approach where each sample may have only a subset of properties labeled, employing masked loss functions to handle missing labels during training.
@@ 3.2.2 Standard MTL with Hard Parameter Sharing
where weights are set inversely proportional to dataset size ($w_{r}=1.0$, $w_{h}=65.0$, $w_{a}=62.0$) to mitigate imbalance.
@@ 4.6.1 Experiment 1: Cost of Data Imbalance
Figure 5 shows the striking result: hardness prediction improves as the data becomes more balanced. With only 1,000 resistivity samples (roughly balanced with 800 hardness samples), hardness $R^{2}=0.830\pm 0.022$. As resistivity samples increase to 52,388, hardness $R^{2}$ degrades to $0.668\pm 0.095$—a 16.3% relative decrease.
@@ 5.1 Why MTL Failed for Regression
5.1.1 Severe Data Imbalance
The most striking characteristic of our dataset is the extreme imbalance: 52,388 resistivity samples versus only 800 hardness samples (65:1 ratio). Despite using inverse-proportion task weights ($w_{h}=65.0$), the optimization dynamics remain dominated by the resistivity task.
During each training batch, the expected number of resistivity samples is approximately $32\times(52388/54028)\approx 31$, while hardness samples appear in only $32\times(800/54028)\approx 0.5$ batches on average. This means most gradient updates are driven entirely by resistivity, with hardness receiving updates only occasionally. The task weights amplify the magnitude of hardness gradients when they occur, but cannot compensate for their infrequency.
Ablation Evidence: Our data quantity ablation (Section 4, Experiment 1) provides direct causal evidence for this hypothesis. When we artificially balanced the data by downsampling resistivity to 1,000 samples, hardness $R^{2}$ improved from 0.668 to 0.830—a 16.3% improvement. This demonstrates that data imbalance alone accounts for the majority of the observed negative transfer.
@@ 5.4.1 Dataset Limitations
5.4.2 Gradient Balancing Methods
We used simple inverse-proportion task weighting, which proved insufficient to overcome severe imbalance. More sophisticated methods like GradNorm [chen2018gradnorm] or uncertainty weighting [kendall2018multi] may yield better results. However, our finding of task independence suggests that even perfect gradient balancing may not enable positive transfer.

### v2

@@ Abstract
Multi-task learning (MTL) has emerged as a promising approach in materials informatics, with the assumption that related material properties share underlying physical principles that can be leveraged to improve prediction accuracy. In this study, we critically examine this assumption by applying MTL to simultaneously predict three properties of metal alloys: electrical resistivity, Vickers hardness, and amorphous-forming ability. Using a union dataset of 54,028 samples from the AI-Hub Korea platform, we compare three approaches: independent single-task models, standard MTL with hard parameter sharing, and structured MTL with learned task relation graphs. Contrary to expectations, our results reveal a striking dichotomy: MTL significantly degrades regression performance (resistivity $R^{2}$: 0.897 $\rightarrow$ 0.844, $p<0.01$; hardness $R^{2}$: 0.832 $\rightarrow$ 0.694, $p<0.01$) while significantly improving classification performance (amorphous F1: 0.703 $\rightarrow$ 0.744, $p<0.05$; recall improved by 17%). Analysis of the learned task relation graphs reveals near-zero inter-task weights ($\sim$0.006), indicating that these alloy properties are fundamentally independent rather than coupled. We attribute the regression degradation to negative transfer caused by severe data imbalance (52,388 resistivity samples vs. 800 hardness samples), where the dominant task overwhelms minority tasks during joint optimization. To address this, we evaluate Deep Imbalanced Regression (DIR) mitigation techniques: PCGrad improves hardness $R^{2}$ by 12.4% (0.761 $\rightarrow$ 0.855), confirming gradient conflicts as the primary bottleneck, while LDS+GradNorm achieves the best overall balance across all tasks. These findings provide empirical guidelines for practitioners: use independent models for precise property regression, leverage MTL with gradient conflict resolution (PCGrad) for minority tasks, and apply LDS+GradNorm when balanced multi-task performance is desired. Our work highlights the importance of validating MTL assumptions before deployment in materials discovery pipelines.
@@ 2.2 Negative Transfer and Task Interference
Chen et al. [chen2018gradnorm] proposed GradNorm to dynamically balance gradient magnitudes across tasks, partially mitigating imbalance effects. Kendall et al. [kendall2018multi] introduced uncertainty-based task weighting. However, these methods address symptoms rather than the fundamental question of task compatibility. A comprehensive survey of negative transfer phenomena is provided by Zhang et al. [zhang2023survey].
2.3 Deep Imbalanced Regression
While class imbalance has been extensively studied in classification, the analogous problem for regression—termed Deep Imbalanced Regression (DIR)—has received attention only recently [yang2021delving]. In DIR, the target variable $y$ is continuous, and the challenge is that certain regions of the target space are underrepresented in training data. This is particularly relevant for materials science, where experimental measurements often cluster around common property values while extreme values are rare.
Yang et al. [yang2021delving] formalized DIR and proposed two key techniques. First, Label Distribution Smoothing (LDS) estimates the effective label density by convolving the empirical distribution with a symmetric kernel:
$\tilde{p}(y)=\int K(y,y^{\prime})p(y^{\prime})dy^{\prime}$ (1)
where $K$ is typically a Gaussian kernel. This smoothed density enables principled reweighting of samples from underrepresented regions. Second, Feature Distribution Smoothing (FDS) applies similar smoothing in feature space to prevent overfitting to majority-region features.
Ren et al. [ren2022balanced] extended this work with Balanced MSE, which reformulates the standard MSE loss to account for label imbalance:
$\mathcal{L}_{BMC}=-\log\mathbb{E}_{y\sim p_{train}(y)}\left[\mathcal{N}(y|f(x),\sigma^{2})\right]$ (2)
This formulation prevents the model from collapsing predictions toward the mode of the training distribution.
These techniques are directly relevant to MTL scenarios with severely imbalanced task sizes. When one task dominates training (as in our 65:1 resistivity-to-hardness ratio), the shared backbone effectively experiences a form of “task-level” imbalanced regression, where the minority task’s signal is overwhelmed by the majority task’s gradients.
@@ 2.5 MTL for Metal Alloys
Debnath et al. [debnath2025overcoming] used MTL to predict yield strength and elongation of HEAs, finding modest improvements. Pei et al. [pei2020machine] predicted hardness and phase stability using composition features. More recently, Kamnis et al. [kamnis2024transformer] demonstrated that transformer-based language models can effectively predict HEA mechanical properties by treating compositions as sequences, achieving strong performance through sequential transfer learning (pre-train then fine-tune) rather than simultaneous MTL. Rao et al. [rao2022machine] showed that active learning approaches, which iteratively select informative samples rather than training on static archives, can discover novel Invar HEAs with high efficiency.
However, most existing studies use small, curated datasets where all samples have all properties labeled—a scenario quite different from the realistic setting where different properties are measured on different sample sets. Furthermore, the question of whether properties from different physical domains (e.g., electronic transport versus mechanical plasticity) can share representations remains largely unexamined.
Our work addresses this realistic scenario using a union dataset approach where each sample may have only a subset of properties labeled, employing masked loss functions to handle missing labels during training. Critically, we examine MTL across physically distinct property types, revealing fundamental limitations of the shared-representation assumption.
@@ 3.2.2 Standard MTL with Hard Parameter Sharing
where weights are set inversely proportional to dataset size ($w_{r}=1.0$, $w_{h}=65.0$, $w_{a}=62.0$) to mitigate imbalance. To address this limitation, we additionally implemented and evaluated several advanced techniques from the Deep Imbalanced Regression (DIR) literature: Label Distribution Smoothing (LDS) [yang2021delving], Balanced MSE [ren2022balanced], Feature Distribution Smoothing (FDS) [yang2021delving], PCGrad [yu2020gradient] for gradient conflict resolution, and GradNorm [chen2018gradnorm] for dynamic task weight balancing. Results of these mitigation experiments are presented in Section 4.
3.2.3 Soft Parameter Sharing MTL
In contrast to hard parameter sharing, soft parameter sharing allows each task to have its own backbone while encouraging similarity through regularization:
$\mathcal{L}_{soft}=\mathcal{L}_{MTL}+\lambda\sum_{i\neq j}\|\theta_{i}-\theta_{j}\|_{2}^{2}$ (4)
where $\theta_{i}$ are task-specific backbone parameters and $\lambda\in\{0.001,0.01,0.1\}$ controls the strength of the similarity constraint. This allows tasks to learn slightly different representations while still benefiting from shared structure.
@@ 3.4 Evaluation Metrics
  • 4.
    Tail metrics: P95/P99 error percentiles, maximum error, and worst-10% MAE for evaluating performance on extreme values
@@ 4.1 Main Performance Comparison
Figure 1: Overall performance comparison between independent and MTL models across all tasks. MTL degrades regression but improves classification.
@@ 4.6.1 Experiment 1: Cost of Data Imbalance
Figure 6 shows the striking result: hardness prediction improves as the data becomes more balanced. With only 1,000 resistivity samples (roughly balanced with 800 hardness samples), hardness $R^{2}=0.830\pm 0.022$. As resistivity samples increase to 52,388, hardness $R^{2}$ degrades to $0.668\pm 0.095$—a 19.5% relative decrease.
@@ 4.6.3 Experiment 3: Transfer Utility Assessment
4.6.4 Experiment 4: Deep Imbalanced Regression Mitigation
Having established that data imbalance and gradient conflicts are the primary drivers of negative transfer, we evaluated whether advanced mitigation techniques from the Deep Imbalanced Regression (DIR) literature could recover performance. We implemented and tested eight configurations combining:
  • 1.
    Loss functions: Standard MSE, Label Distribution Smoothing (LDS) [yang2021delving], Balanced MSE [ren2022balanced]
  • 2.
    Gradient optimization: Standard backpropagation, PCGrad [yu2020gradient], GradNorm [chen2018gradnorm]
Table 6 presents the comprehensive comparison:
Table 6: DIR mitigation experiment results. Best results for each metric are bolded. $\Delta$ indicates improvement over baseline MSE.
Method Resistivity $R^{2}$ $\Delta$ Hardness $R^{2}$ $\Delta$ Amorphous F1 $\Delta$
Baseline MSE 0.874 0.761 0.692
LDS Only 0.869 -0.6% 0.767 +0.8% 0.729 +5.3%
Balanced MSE 0.869 -0.6% 0.767 +0.8% 0.729 +5.3%
GradNorm Only 0.873 -0.1% 0.786 +3.3% 0.744 +7.5%
PCGrad Only 0.852 -2.5% 0.855 +12.4% 0.701 +1.3%
Balanced MSE + PCGrad 0.850 -2.8% 0.827 +8.7% 0.676 -2.3%
LDS + PCGrad 0.850 -2.8% 0.827 +8.7% 0.676 -2.3%
LDS + GradNorm 0.894 +2.4% 0.789 +3.7% 0.750 +8.4%
Key Finding 1: PCGrad confirms gradient conflict hypothesis. PCGrad alone achieved the largest improvement for hardness prediction: $R^{2}$ improved from 0.761 to 0.855 (+12.4%). This dramatic improvement—larger than any other intervention—provides strong empirical evidence that gradient conflicts between tasks were indeed harming the minority task. By projecting conflicting gradient components, PCGrad prevents the majority task (resistivity) from overwriting gradients beneficial to hardness prediction.
Key Finding 2: LDS+GradNorm achieves best overall balance. While PCGrad excels at protecting the minority task, it comes at a cost to the majority task (resistivity $R^{2}$ dropped from 0.874 to 0.852). In contrast, LDS+GradNorm achieves improvements across all tasks simultaneously: resistivity $R^{2}$ improved to 0.894 (+2.4%), hardness $R^{2}$ to 0.789 (+3.7%), and amorphous F1 to 0.750 (+8.4%). GradNorm’s dynamic task weight adjustment prevents any single task from dominating, while LDS smooths the label distribution to improve minority sample handling.
Key Finding 3: No synergy between LDS and PCGrad. Combining LDS with PCGrad (the “hero run”) yielded worse results than PCGrad alone for hardness ($R^{2}=0.827$ vs. 0.855). The synergy score is negative:
$\text{Synergy}=R^{2}_{\text{LDS+PCGrad}}-\max(R^{2}_{\text{LDS}},R^{2}_{\text{PCGrad}})=0.827-0.855=-0.028$ (8)
This suggests that these techniques may be addressing overlapping aspects of the imbalance problem, and combining them introduces interference rather than complementary benefits.
Conclusion: DIR mitigation techniques can substantially improve MTL performance, but the optimal choice is task-dependent. For maximizing minority-task performance, PCGrad alone is optimal. For balanced multi-task performance, LDS+GradNorm is preferred.
4.6.5 Experiment 5: Feature Distribution Smoothing (FDS)
Feature Distribution Smoothing (FDS) [yang2021delving] estimates feature-space density using kernel density estimation and upweights samples in sparse regions. We implemented FDS with Gaussian kernels ($\sigma=2$) to address the uneven distribution of samples in feature space that accompanies our label imbalance.
Table 7 shows the results:
Table 7: Feature Distribution Smoothing results. FDS shows no improvement over baseline.
Method Resistivity $R^{2}$ Hardness $R^{2}$ Amorphous F1
Baseline MSE 0.874 0.761 0.692
FDS Only 0.874 0.761 0.692
LDS + FDS 0.874 0.761 0.692
Key Finding: FDS provides no measurable improvement. Unlike natural image datasets where FDS was originally developed, our alloy composition features are already well-distributed in the compositional simplex. The feature imbalance addressed by FDS is distinct from the label imbalance that causes negative transfer in our setting.
4.6.6 Experiment 6: Soft Parameter Sharing
Hard parameter sharing (shared backbone) forces all tasks to learn identical representations, which may not be optimal when tasks have different requirements. We implemented soft parameter sharing with separate task-specific backbones regularized by L2 penalties encouraging similarity:
$\mathcal{L}_{\text{soft}}=\sum_{i\neq j}\lambda\|\theta_{i}-\theta_{j}\|_{2}^{2}$ (9)
where $\theta_{i}$ are task-specific backbone parameters and $\lambda$ controls the similarity pressure.
Table 8 compares hard and soft parameter sharing across different $\lambda$ values:
Table 8: Soft parameter sharing results. $\lambda=0.001$ improves hardness and classification while maintaining resistivity performance.
Config Resistivity $R^{2}$ Hardness $R^{2}$ Amorphous F1 Backbone Sim.
Hard Sharing 0.875 0.735 0.714 1.000
Soft $\lambda=0.001$ 0.868 0.764 0.744 0.996
Soft $\lambda=0.01$ 0.852 0.692 0.750 0.999
Soft $\lambda=0.1$ 0.849 0.704 0.774 0.999
Key Finding 1: Soft sharing improves minority task performance. With $\lambda=0.001$, hardness $R^{2}$ improves from 0.735 to 0.764 (+3.9%) and amorphous F1 improves from 0.714 to 0.744 (+4.2%), at a small cost to resistivity (-0.8%). The learned backbone similarity remains high (99.6%), indicating that task-specific deviations are subtle but beneficial.
Key Finding 2: Higher $\lambda$ forces convergence but hurts performance. As $\lambda$ increases, backbones converge to near-identical representations (¿99.9% similarity), approaching hard sharing. Counter-intuitively, this degrades hardness performance ($R^{2}=0.692$ at $\lambda=0.01$), suggesting that the minority task benefits from having slightly different representations.
Key Finding 3: Soft sharing with tail metrics analysis. Table 9 shows worst-case metrics for soft parameter sharing:
Table 9: Tail metrics for soft parameter sharing ($\lambda=0.001$). Soft sharing reduces worst-case errors for hardness.
Task Max Error P95 Error P99 Error Worst-10% MAE
Hard Sharing
Resistivity 2.19 0.60 1.06 0.69
Hardness 3.11 1.00 1.80 1.22
Soft $\lambda=0.001$
Resistivity 2.07 0.63 1.06 0.71
Hardness 1.95 1.05 1.39 1.12
Soft parameter sharing reduces the maximum hardness prediction error from 3.11 to 1.95 (37% reduction), indicating improved robustness for extreme values.
Conclusion: Soft parameter sharing provides a complementary approach to gradient-based methods. While PCGrad achieves higher overall hardness $R^{2}$ (0.855 vs. 0.764), soft sharing may be preferred when tail performance is critical or when gradient manipulation is undesirable.
@@ 5.1 Why MTL Failed for Regression
5.1.1 Deep Imbalanced Regression Analysis
Our observations align closely with the Deep Imbalanced Regression (DIR) framework recently formalized by Yang et al. [yang2021delving]. The most striking characteristic of our dataset is the extreme imbalance: 52,388 resistivity samples versus only 800 hardness samples (65:1 ratio). This manifests as two distinct types of imbalance:
  1. 1.
    Task-level imbalance: The sheer difference in sample counts means the majority task dominates gradient updates.
  2. 2.
    Target-value imbalance: Within the hardness dataset itself, certain property ranges may be underrepresented, compounding the difficulty of accurate regression.
The standard MSE loss function is fundamentally ill-suited for this scenario:
$\mathcal{L}_{total}=\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}$ (10)
In the union dataset, the vast majority of terms come from the resistivity task. Even with task weighting, the gradient density remains skewed—the model updates its shared parameters tens of thousands of times to satisfy the resistivity manifold for every few hundred updates toward hardness.
Why Inverse Weighting is Insufficient. Despite using inverse-proportion task weights ($w_{h}=65.0$), the optimization dynamics remained dominated by the resistivity task. This failure can be attributed to several factors identified in the DIR literature:
  • 1.
[…]
  • 2.
    Feature shift: The backbone learns representations optimized for resistivity prediction. When hardness gradients occasionally propagate, they attempt to reshape these features, but the overwhelming “gravitational pull” of the majority data quickly pulls the representations back.
  • 3.
    Gradient starvation: The minority task effectively starves during training, receiving insufficient gradient signal to shape the shared representation [yu2020gradient].
DIR Mitigation Results. To address these limitations, we implemented and evaluated several techniques from the DIR literature (see Section 4, Experiment 4). The results reveal important insights about gradient conflict resolution:
PCGrad [yu2020gradient] achieved the most dramatic improvement for the minority task: hardness $R^{2}$ improved from 0.761 to 0.855 (+12.4%). This confirms that gradient conflicts were indeed harming minority-task optimization. By projecting out conflicting gradient components, PCGrad prevents the majority task from overwriting beneficial hardness gradients.
GradNorm [chen2018gradnorm] combined with Label Distribution Smoothing (LDS) [yang2021delving] achieved the best overall balance, improving all three tasks simultaneously: resistivity $R^{2}=0.894$ (+2.4%), hardness $R^{2}=0.789$ (+3.7%), and amorphous F1 = 0.750 (+8.4%). Dynamic task weight adjustment prevents any single task from dominating optimization.
Interestingly, combining LDS with PCGrad showed negative synergy (0.028-0.028), with hardness $R^{2}$ dropping from 0.855 to 0.827. This suggests that these techniques address overlapping aspects of the imbalance problem and may interfere when combined.
Ablation Evidence: Our data quantity ablation (Section 4, Experiment 1) provides direct causal evidence for the DIR hypothesis. When we artificially balanced the data by downsampling resistivity to 1,000 samples, hardness $R^{2}$ improved from 0.668 to 0.830—an absolute improvement of 0.162. This demonstrates that data imbalance alone accounts for the majority of the observed negative transfer, consistent with the DIR framework’s predictions.
@@ 5.1.3 Representation Conflict
5.1.4 Functional Form Mismatch
A deeper explanation for task independence emerges from considering the functional forms that govern different material properties. Recent work in symbolic regression for materials science [wang2019symbolic] has revealed that different properties follow qualitatively different mathematical dependencies on composition.
Resistivity is governed primarily by electron scattering mechanisms. For solid solutions, the classical Nordheim rule [nordheim1931resistivity] gives:
$\rho\approx\rho_{0}+\sum_{i}k_{i}c_{i}(1-c_{i})$ (11)
where $c_{i}$ is the concentration of solute $i$ and $k_{i}$ captures the scattering strength. This is a polynomial form, symmetric about $c_{i}=0.5$.
Hardness and yield strength, by contrast, follow power-law dependencies related to lattice distortion and solid-solution strengthening [toda2015modelling, sengupta2021review]:
$\sigma_{y}\approx A+B\delta^{2/3}c^{1/2}$ (12)
where $\delta$ is the atomic size mismatch parameter and $c$ is the solute concentration. The fractional exponents ($2/3$, $1/2$) arise from dislocation-solute interaction physics.
A neural network backbone fundamentally performs a series of linear and nonlinear transformations to create a latent representation. If Task A requires a representation proportional to $c_{i}(1-c_{i})$ (parabolic) and Task B requires $c^{1/2}$ (square-root), the backbone faces a fundamental tension. With balanced data, the model might learn a “super-representation” containing both feature types. However, with our 65:1 imbalance, the backbone learns only the polynomial form optimized for resistivity. When the hardness prediction head receives these features, it attempts to fit a power-law relationship using polynomial features—a mathematical mismatch that results in systematic prediction errors.
This functional form mismatch explains why our transfer learning experiment (Section 4, Experiment 3) showed no benefit from pre-training on resistivity: the learned representations encode the wrong inductive bias for hardness prediction.
@@ 5.3.3 Recommendation 3: Validate MTL Assumptions Before Deployment
5.4 Toward a Materials Taskonomy
Our findings suggest the need for a systematic “taskonomy” of material properties, analogous to that developed by Zamir et al. [zamir2018taskonomy] for computer vision tasks. Rather than assuming all material properties are related because they derive from atomic-level physics, we should empirically map which properties share sufficient representational structure to benefit from joint learning.
Based on our analysis and the underlying physics, we propose organizing alloy properties into clusters by physical mechanism:
Table 10: Proposed materials property clusters based on underlying physics. MTL is expected to yield positive transfer within clusters but negative transfer across clusters.
Cluster Properties Governing Physics
Electronic Resistivity, Thermal Conductivity, Seebeck Electron transport/scattering
Mechanical Hardness, Yield Strength, UTS Dislocation mechanics
Thermodynamic Formation Enthalpy, Phase Stability Free energy landscapes
Key insight: The negative transfer observed in our study occurred precisely because we attempted to bridge the Electronic cluster (resistivity) and the Mechanical cluster (hardness) using a shared representation dominated by Electronic-task data. Properties within the same cluster—such as hardness and yield strength—share common functional dependencies and are likely to exhibit positive transfer. Properties across clusters follow different mathematical forms and may actively interfere when forced to share representations.
This taskonomy framework has practical implications: when constructing MTL models for materials discovery, practitioners should preferentially group properties within the same physical cluster. Cross-cluster MTL should be approached with caution and requires explicit validation against independent baselines.
@@ 5.5.1 Dataset Limitations
5.5.2 Effectiveness of DIR Mitigation Techniques
Our comprehensive evaluation of DIR techniques (Section 4, Experiment 4) reveals nuanced findings about when each approach is effective:
  • 1.
    PCGrad [yu2020gradient]: Highly effective for minority-task protection. The +12.4% improvement on hardness $R^{2}$ confirms that gradient conflicts were the primary bottleneck for this task. However, it comes at a cost to majority-task performance (-2.5% on resistivity).
  • 2.
    GradNorm [chen2018gradnorm]: Provides balanced improvements across all tasks through dynamic weight adjustment. Combined with LDS, achieves the best overall performance profile.
  • 3.
    LDS and Balanced MSE [yang2021delving, ren2022balanced]: Modest improvements for classification (+5.3% F1) but limited benefit for regression tasks when used alone.
  • 4.
    Negative synergy: Combining LDS with PCGrad degrades performance compared to PCGrad alone, suggesting these techniques address overlapping mechanisms and may interfere.
These results refine our understanding of task independence. While properties from different physical clusters (electronic vs. mechanical) cannot benefit from shared representations, gradient conflict resolution can still substantially improve minority-task performance by preventing destructive interference during optimization. The functional form mismatch between electronic and mechanical properties remains a structural limitation, but its impact can be mitigated through careful gradient management.